# Social Genome: Grounded Social Reasoning Abilities of Multimodal Models

### v2

@@ 5 Conclusion
This paper introduces Social Genome, the first benchmark for fine-grained, grounded social reasoning abilities of multimodal models. The reasoning traces contributed by Social Genome include references to multimodal cues and external knowledge concepts that humans find useful when performing social inferences. We define metrics to assess semantic and structural aspects of reasoning traces, and we contribute novel insights regarding gaps and opportunities to improve the grounded social reasoning capabilities of multimodal models.
Future AI systems interacting with humans in face-to-face, embodied contexts must have the ability to reason about social interactions and ground reasoning in concrete multimodal evidence and external knowledge concepts. Social Genome serves as a first step towards studying and advancing this form of reasoning in AI systems.
@@ 6 Limitations
Video Contexts
The videos in Social Genome each have a length of 1 minute. We believe that findings from Social Genome regarding model social reasoning abilities about interactions in these 1-minute clips motivate future work to curate longer-form video benchmarks. Social interactions during hour-long videos Zou et al. (2025), for example, would contain more complex social dynamics and longer-form temporal dependencies across humans. In addition, all videos in Social Genome have social interactions in English, and all annotators were required to be proficient in English. As social norms can differ across cultures and languages Hershcovich et al. (2022), we believe future research in multimodal social reasoning would benefit from a larger-scale curation of multimodal interaction data across cultural and linguistic contexts.
@@ 7 Ethics
Social reasoning abilities are essential for future AI systems to effectively work with and alongside humans in real-world settings. Social Genome has the potential to support the research community in studying and advancing these capabilities in AI systems. We envision AI systems using social reasoning to enhance human autonomy, health, and well-being – for example, a healthcare agent reasoning about a patient’s nonverbal behaviors, in order to help doctors better support them. However, these technologies exist within a broader societal context, with multiple stakeholders and potential risks in areas such as privacy, surveillance, and manipulation. We support broader academic research and policy efforts to mitigate against misuse and potential harms of socially-intelligent AI.
Acknowledgements
@@ A.1 Social Genome Data Sourcing
All videos in Social Genome were sourced from publicly-available Social-IQ 2.0 modeling challenge Wilf et al. (2023). We obtained permission from the authors to access the original test set videos and question-answer tuples from this dataset (275 videos and 1514 QA Tuples). Our use of these videos aligns with Social-IQ 2.0 repository’s MIT license and intended research purpose. We chose to use these test set videos as candidate seed videos to build Social Genome because the answers to questions posed about these videos have not been released online, reducing chances of benchmark contamination Xu et al. (2024). At the current stage, we plan to avoid benchmark contamination in Social Genome by not uploading human social reasoning trace annotations online and using them for research purposes.
@@
A.2 Social Genome Annotation
@@
A.3 Social Genome Validation
@@
A.4 Social Genome Human Accuracy
@@ C.2 Model Generation Notes
We note observations here on model generations. VideoChat generations for $k\in{0,1,2,4}$ produced full sentences explaining reasoning traces, however the generation quality eroded for $k\in{8,16}$. Several samples from these settings of $k$ were repeated words and short phrases (e.g.,“ the the the...and and and and"). Similarly, VideoChat-GPT generations for $k\in{0,1,2,4}$ produced full sentences explaining reasoning, however the generation quality eroded for $k\in{8,16}$. Samples from these settings of $k$ were repeated short words and letters (e.g.,“or or or, or" and “B ( ( ( ( ( B"). LLaVA-Video generations for $k\in{0,1,2,4,8,16}$ produced full sentences explaining reasoning, however the generations as $k$ increased in $k\in{4,8,16}$ began to answer fewer questions. LLaVA-Video-Only answered questions, but did not generate reasoning traces; this model was discussed in Section 4 solely for the social inference accuracy metric. These model generation challenges were not observed for LongVA or Gemini-1.5. GPT-4o initially generated “I’m sorry, I can’t assist with that." as one of the reasoning trace steps for several questions, before answering the question.
@@ D.2 Does Chain-of-Thought Prompting Elicit Multimodal Social Reasoning?
Appendix E Human Evaluation Criteria

### v3

@@ 5 Conclusion
We introduce Social Genome, the first benchmark for fine-grained, grounded social reasoning abilities of multimodal models. Reasoning traces contributed by Social Genome include multimodal cues and external knowledge concepts that humans find useful when performing social inferences. We define metrics to assess semantic and structural aspects of reasoning traces and contribute novel insights regarding gaps and opportunities to improve the grounded social reasoning capabilities of multimodal models. Future AI systems reasoning about social interactions must be able to ground reasoning in concrete multimodal evidence and external knowledge concepts. Social Genome serves as a first step towards studying and advancing this form of reasoning in AI systems.
@@ 6 Limitations
Video Lengths
The videos in Social Genome each have a length of $\sim$1 minute, consistent with the lengths of existing video understanding benchmarks (e.g., Social-IQ 1.0 and Social-IQ 2.0 have 1-minute samples Wilf et al. (2023); Zadeh et al. (2019), TVQA has 1.3 minute samples Lei et al. (2018), and MEmoR has 30 second samples Shen et al. (2020)). Social reasoning regularly occurs in micro-social and shorter-term contexts; humans make split-second inferences about emotions Nook et al. (2015), social behaviors and gestures Beattie and Aboudan (1994), and personality Lin et al. (2021), among other social phenomena. The interactions in Social Genome videos contain rich, nuanced social signals and multimodal behavioral dynamics that require social reasoning to interpret. Our paper demonstrates that current state-of-the-art models struggle to interpret 1-minute social interactions. The length of our videos is not, in itself, a technical limitation of our research; however, we would like to motivate the need for community-driven curation of longer-form social interaction datasets in future years.
Scope of the Study
Videos in Social Genome have interactions in English, and annotators were required to be proficient in English. The study was scoped within these constraints, consistent with prior multimodal video understanding tasks (Table 1). A multilingual and multicultural data collection was not within the scope of this research. Our paper motivates future research in multimodal social reasoning that includes a community-driven curation of interaction data across sociocultural contexts.
@@ 7 Ethics
Bias Considerations
Annotators in the original Social-IQ 2.0 dataset, from which we sourced seed videos, used terms such as "man" in alignment with annotator perception of gender. In Social Genome we did not frame judgements about gender identity of individuals based on these annotations. In Social Genome, 45% of samples refer to women, and 17% make no reference to gender. Samples involving women do not have reasoning traces that refer more frequently to emotion words ($r$ = 0.033). We find no substantial difference in model performance across gender; for example, Gemini-1.5-Flash social inference accuracy is 74.9% for samples solely involving women, 73.4% for sampling solely involving men, 73.6% for samples referring to multiple genders, and 77% for samples that do not specify gender.
Environmental Statement
Experiments used a single A100 GPU, a carbon footprint of 1.24 kgCO2e, and an energy consumption of 3.72 kWh [^3] http://calculator.green-algorithms.org.
@@
Social reasoning abilities are essential for future AI systems to effectively work with and alongside humans. Social Genome has the potential to support the research community in studying and advancing these capabilities in AI systems. We envision AI systems using social reasoning to enhance human autonomy, health, and well-being. However, these technologies exist with potential risks in amplifying toxicity Zhou et al. (2023), surveillance, and manipulation. We support broader research and policy efforts to mitigate against misuse and potential harms of socially-intelligent AI.
Acknowledgments
@@ A.1 Social Genome Data Sourcing
All videos in Social Genome were sourced from publicly-available Social-IQ 2.0 modeling challenge Wilf et al. (2023). We obtained permission from the authors to access the original test set videos and question-answer tuples from this dataset (275 videos and 1514 QA Tuples). Our use of these videos aligns with Social-IQ 2.0 repository’s MIT license and intended research purpose. We chose to use these test set videos as candidate seed videos to build Social Genome because the answers to questions posed about these videos have not been released online, reducing chances of benchmark contamination Xu et al. (2024). At the current stage, we plan to avoid benchmark contamination in Social Genome by not uploading human social reasoning trace annotations online, by using them for research purposes, and by maintaining a leaderboard for the community.
@@
A.2 Social Genome  Entities
Across the 1486 samples and 11,253 entities (people, objects, concepts) mentioned in Social Genome, we visualize the distribution of non-human entities (objects, concepts) in Figure 5, with an average of 7.6 unique entities referenced per human-annotated reasoning trace. Figure 6 visualizes entities most frequently mentioned by human annotators (not including words such as "man" and "woman" repeated from question statements). As seen in Figure 6, human annotators constructing social reasoning traces focused on multimodal aspects of social interactions, with evidence spanning vocal (e.g., "tone", "voice"), visual (e.g., "eyes", "head", "body"), and verbal (e.g, "words", "conversation") cues. These observations support the perspective that interpreting and reasoning about real-world social interactions requires an integration of multimodal information.
Figure 5: Distribution of entity counts in Social Genome samples.
Figure 6: Top entity counts mentioned by human annotators in Social Genome reasoning traces.
Human annotators mentioned an average of 2.23 emotions per reasoning trace. Further analysis of the emotions referenced indicated that these emotions spanned a diverse range. Emotions frequently mentioned were the following: happy, serious, surprised, excited, nervous, calm, confused, angry, annoyed, comfortable, sarcastic, and positive.
A.3 Social Genome Annotation
@@
A.4 Social Genome Validation
@@
A.5 Social Genome Human Accuracy
@@ C.2 Model Generation Notes
We note observations here on model generations. VideoChat generations for $k\in{0,1,2,4}$ produced full sentences explaining reasoning traces, however the generation quality eroded for $k\in{8,16}$. Several samples from these settings of $k$ were repeated words and short phrases (e.g.,“ the the the...and and and and").
Similarly, VideoChat-GPT generations for $k\in{0,1,2,4}$ produced full sentences explaining reasoning, however the generation quality eroded for $k\in{8,16}$. Samples from these settings of $k$ were repeated short words and letters (e.g.,“or or or, or" and “B ( ( ( ( ( B").
LLaVA-Video generations for $k\in{0,1,2,4,8,16}$ produced full sentences explaining reasoning, however the generations as $k$ increased in $k\in{4,8,16}$ began to answer fewer questions. LLaVA-Video-Only answered questions, but did not generate reasoning traces; this model was discussed in Section 4 solely for the social inference accuracy metric.
These model generation challenges were not observed for LongVA or Gemini-1.5. GPT-4o initially generated “I’m sorry, I can’t assist with that." as one of the reasoning trace steps for several questions, before answering the question.
@@ D.2 Does Chain-of-Thought Prompting Elicit Multimodal Social Reasoning?
Appendix E Human Evaluation Details
@@
Our automated metrics yield similar model rankings to human evaluation and strong correlation to human judgements, supporting the reliability of these metrics as proxies for reasoning trace quality. Gemini-1.5-Flash achieves the highest score in human judgements for referencing “Low-Level” behavioral cues and achieves the highest score in automated metrics such as Accuracy, overall semantic similarity (Similarity-Trace), and step-level semantic similarity (Similarity-Num Steps). Similarly, GPT-4o and LLaVA-Video, followed closely by LongVA, score higher than Video-ChatGPT and VideoChat2 on both automated metrics and human judgments. This rank-based alignment indicates that automated proxy metrics can capture reasoning quality signals that human evaluators identified.
These trends observed from rank-based alignment are supported by correlation analyses between model metrics and human judgments. For Gemini-1.5-Flash and LLaVA-Video (highest-performing closed-source model and open-source model), the Similarity-Trace and Similarity-Step metrics exhibit $R^{2}$ values of 0.75 and 0.61, respectively, demonstrating strong alignment with human judgments of how effectively models ground reasoning in low-level behavioral cues. Strong correlations indicate that (1) these automated metrics can be viewed as effective predictors of low-level social reasoning quality and (2) there is room for future community research into automated evaluation of social reasoning quality. Social Genome contributes a new benchmark and findings for the community to study this direction.